OpenStudy (anonymous):

Can you solve this system. {x^2+5x+5=y x+1=y Can you graph both equations and label the vertex of parabola, x-intercept and y-intercept for both equations?

OpenStudy (campbell_st):

well since they both equal y you can equate them so \[x^2 + 5x + 5 = x + 1\] just solve for x

OpenStudy (anonymous):

I really don't know how to do that?? @campbell_st

OpenStudy (campbell_st):

ok... start by subtracting x from both sides and then 1 you get \[x^2 + 4x + 4 = 0\] can you solve that..?

OpenStudy (anonymous):

im feeling clueless